The activation energy for the decomposition of hydrogen iodide is found to be 180 kJ mol–1 and the rate constant for this reaction is reported to be 0.24 mol–1 L min–1at 540 °C. What is k for this reaction at 30 °C?
Answer #1


ln(k2/k1) = Ea/R[1/T1 - 1/T2]

K1 = ?           T1 = 30 c = 303 k

K2 = 0.24        T2 = 540 c = 813 k

Ea = 180 kj/mol , R = 8.314 j.k-1.mol-1

ln(0.24/k1) = (180*10^3/8.314)((1/303)-(1/813))

K1 = rate constant at 30 c = 8.2*10^-21